OpenStudy (lgg23):

Evaluate the indefinite integral of e^(4x)sin(2x) dx

OpenStudy (bahrom7893):

integration by parts.. ughh longgggg

OpenStudy (bahrom7893):

http://www.wolframalpha.com/input/?i=integrate+%28e^%284x%29sin%282x%29+dx%29 click on show steps

OpenStudy (ash2326):

We have \[\int e^{4x} \sin 2x dx\] we'll use integration by parts \[\int uv=u\int v- \int (v' \int u)\] here \[I=\int e^{4x} \sin 2x dx\] here \[v=e^4x\] \[u=\sin2 x dx\] we now integrate I using by parts \[I= e^{4x}\int\sin 2x dx- \int 4e^{4x }(\int \sin 2x dx) dx\] we have \[ I = -e^{4x} \cos 2x /2+ \int 4 e^{4x} \cos 2x /2dx\] we have now \[I=-e^{4x} \cos 2x /2+ \int 2e^{4x} \cos 2x dx\] let's integrate the second term using by parts \[I=-e^{4x} \cos 2x /2+ 2e^{4x}\int \cos 2x dx-\int \frac{d}{dx}2e^{4x} (\int \cos 2x dx) dx\] we get now \[I=-e^{4x} \cos 2x /2+ 2e^{4x}\sin 2x /2-\int 8 e^{4x} \sin 2x /2 dx\] we know that \[I=\int e^{4x} \sin 2x dx\] the second term is -4I so we get \[I=-e^{4x} \cos 2x /2+ 2e^{4x}\sin 2x /2-4I\] so we get \[5I=-e^{4x} \cos 2x /2+ 2e^{4x}\sin 2x /2\] we get finally \[I=\frac{e^{4x}}{5}(-\cos 2x /2+ \sin 2x )\]

OpenStudy (ash2326):

forgot to add constant , add C with the answer